OpenStudy (anonymous):

Can 3√2x be simplified? I know x^2 can be simplified as x*x but is there a way to simplify 2x?

hartnn (hartnn):

\(\sqrt{2x}=\sqrt{2}\sqrt{x}\) but i wouldn't call that a simplification :P

hartnn (hartnn):

so, \(3\sqrt{2x}\) is in most simplified form

OpenStudy (anonymous):

oh okay, just making sure. i was trying to simplify one at a time to: 3√2x + x√8x - 5√18x thats why i was curios lol. but thank you!

hartnn (hartnn):

you need help simplifying 3√2x + x√8x - 5√18x ?

hartnn (hartnn):

or you are trying first, if you need help you can ask me after u've tried...

OpenStudy (anonymous):

I tried to simplify it and got 3√2x + 2√2x - 15√2x And then I got -15√2x but I'm not sure if that's right or not.

hartnn (hartnn):

did you just change the question ? :P

hartnn (hartnn):

or was x√8x a typo and its just √8x

hartnn (hartnn):

i assume its √8x 3√2x + 2√2x - 15√2x √2x [3+2-15] = √2x [.....?...]

OpenStudy (anonymous):

oh i didnt mean to change the question! its x√8x

hartnn (hartnn):

:O then, we have 3√2x + 2x√2x - 15√2x √2x [3+2x-15] = √2x [.....?...]

OpenStudy (anonymous):

√2x - 12 ? Would that go before the radical?

hartnn (hartnn):

how you got √2x - 12 ? 3+2x-15 = 2x-12 so, √2x [3+2x-15] = √2x [2x-12] thats it!

OpenStudy (anonymous):

oh okay, i just thought because it was after the radical. thank you!

hartnn (hartnn):

welcome ^_^